# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v3

@@ III Results
The temporal evolution of the system reveals a spontaneous transition from microscopic chaos to macroscopic order (Figure 1). Initially ($t<10s$), the system is dominated by small-scale phase defects that nucleate and annihilate rapidly. However, as the system evolves, surviving defects begin to cluster and merge, and by $t=200s$, the flow is dominated by a sloshing, large-scale Onsager dipole.
@@
To elucidate the mechanism driving this condensation, we analyzed the kinetic pathways of the topological defects (Figure 2b), the decay of the defect number density $N_{d}(t)$. Here, adherence to the theoretical $t^{-0.75}$ scaling for vortex packing [36], significantly slower than the diffusive $t^{-1}$ rate, confirms that vortex interaction is governed by long-range strain fields (vortex thinning [37, 38]) rather than random-walk annihilation [39, 40].
@@
The directionality of the energy transfer is established by the spectral energy flux $\Pi(k)$ (Figure 2c), where the distinct negative plateau in the spectral flux ($\Pi(k)<0$) rigorously proves the inverse directionality of the energy cascade; energy injected at the microscale by $\omega_{i}$ is transported upscale to the box size, $\sim 2L$, creating the Onsager dipole.
@@
A central finding of this work is the resolution of the spectral dichotomy observed in active matter, namely the coexistence of steep, dissipative spectra with large-scale coherent structures [12, 41, 42]. Figure 2d) shows that raw particle velocity field exhibits a steep spectral slope $E_{raw}(k)\propto k^{-3}$, consistent with dissipative enstrophy [30]. This reflects the singular geometry of the topological defects: the raw agents are kinematically pinned to the sharp cusps of the phase field ($\nabla\phi\sim 1/r$), and thus their spectrum measures the packing of vorticity filaments [43], exactly analogous to forced two-dimensional turbulence [33].
However, when we apply the RFE operator to the ensemble, filtering out the microscopic singularities by averaging over the interaction scale $\tau$, a “hidden” Kolmogorov scaling is revealed: The RFE spectrum $E_{RFE}(k)$ (the blue spectrum in Figure 2d) tracks the classical $k^{-5/3}$ slope characteristic of 2D inviscid fluid turbulence, confirming our hypothesis that the chiral active matter (the oscillatory particles) acts as a dissipative kinetic layer that drives, through inverse cascade, a superfluid (the RFE field) governed by the conservative Kirchhoff-Onsager Hamiltonian.
@@
Finally, we demonstrate that the formation of the dipole is a thermodynamic inevitability. Figure 2d) maps the system’s trajectory in the energy-order phase plane ($E$ vs. $P$). The system climbs monotonically away from the disordered origin (random gas), increasing both its macroscopic kinetic energy and its dipole order parameter $P$. In the context of the point-vortex Hamiltonian, this trajectory corresponds to an ascent up the entropy curve into the regime of negative absolute temperature [19]. The intrinsic disorder ($\omega_{i}$, or rather, the distribution $\Delta\omega$) continuously injects enstrophy to maximize the entropy of the renormalised fluid, forcing the spectral condensation of vorticity as the only accessible maximum-entropy state within the bounded domain. This unifies the phenomenology of chiral active swarms with the rigorous statistical mechanics of classical inviscid fluids.
@@
Figure 3: Active vortex glass made of particles whose raw distribution features an inverse cascade ($\Pi(k)<0$), but whose RFE field features a forward-cascade ($\Pi(k)>0$). The defect count falls off drastically and flattens; defect merger is arrested and the ensemble settles in a frozen defect lattice. Top row shows the particle point-cloud while bottom row shows topological charge, while the columns correspond to seven temporal snapshots.
@@ IV Discussion
Our results resolve a spectral paradox [44, 45, 4, 46] by demonstrating that chiral active matter may not be governed by a single physical regime, but by a scale-dependent duality. At the microscopic scale ($k>k_{c}$), the system is dominated by phase singularities, with a steep spectral slope ($E_{raw}\propto k^{-3}$) [30, 31, 32, 33]; this is not a signature of dissipation, but of enstrophy injection, the active stretching and packing of vorticity filaments [37, 38]. This layer acts as a thermodynamic furnace that converts intrinsic chemical energy (represented by $\omega_{i}$) into topological charge.
@@
I. Global synchronisation $\Delta\omega\to 0$ $\nabla\phi\to 0$. Static clumps. $\Pi_{Raw}>0,\;\Pi_{RFE}>0$
II. Active Vortex Glass Low / narrow Frustration; no merger. Defect solid. $\Pi_{Raw}<0,\;\Pi_{RFE}>0$
III. Onsager Condensate High / broad Vortex merger via inverse cascade. Onsager dipole. $\Pi_{Raw}<0,\;\Pi_{RFE}<0$
At macroscopic scales ($k<k_{c}$), the system reveals a hidden renormalised fluid. By filtering out the singular defects via the RFE operator (Eqs. 45), we recover the Kolmogorov-$k^{-5/3}$ inverse cascade. This implies that the collective motion of the defects generates a “phantom inertia,” an effective advective term, $(\mathbf{v}\cdot\nabla)\mathbf{v}$ (see Appendix B), arising from the effective, active stress tensor, which then drives the system according to the laws of inviscid hydrodynamics, despite the overdamped nature of the individual agents.
@@
Crucially, the RFE framework suggests a general resolution for the non-universal spectra frequently reported in biological and active matter experiments [9]. We propose that the deviations from Kolmogorov scaling observed in bacterial swarms, active nematics, and cytoskeletal gels [1, 9, 10, 7, 3, 47] may be artifacts of analyzing the raw particle fields rather than the renormalised hydrodynamic modes. By applying the RFE operator (Eqs. 45), essentially a low-pass filter, it may be possible to unveil hidden inertial-like scaling laws in a broad class of overdamped systems.
IV.1 Topological gas dynamics
While the RFE analysis confirms the existence of an underlying hydrodynamic skeleton ($k^{-5/3}$), the realization of the Onsager dipole depends strictly on the system’s ability to overcome topological energy barriers (see, e.g., Ref. [18]). As alluded to by Ref. [17], we find that the intrinsic frequency dispersion $\Delta\omega$ acts as an enstrophy injection rate for the topological gas, determining whether the system freezes, jams, or flows.
Figure 3 illustrates the foregoing by presenting the evolution of a simulation that used a narrower, lower-frequency dispersion $\Delta\omega$ (see Figure 4 in the Appendix A); the raw particle distribution subsequently drove an inverse cascade, but the renormalised fluid did not, and the active bath was unable to transfer energy to wavenumbers $k<k_{c}$, resulting in an arrested toplogical defect glass state. While the Onsager fluid exhibits continuous, power-law coarsening ($N\propto t^{-0.75}$), the vortex glass state exhibits a rapid localized quench ($N\propto t^{-2}$) followed by a kinetic arrest ($N\rightarrow\text{const.}$), yielding a complete absence of topological sorting, but the clear presence of quantized loop currents of phase information [27]; we tentatively note that this may constitute a novel state of matter.
Conversely, the Onsager dipole state, which represents the hydrodynamic limit, is in a marginal synchronisation regime, meaning that the synchronising force is in fact negligible for the instantaneous angular velocity, but maintains a non-zero order parameter $R$ (facilitating flow). This is empirically confirmed by observing that the dynamics in Figure 1 effectively never stop. If $R\rightarrow 1$, the hydrodynamic limit is never reached, and the ensemble experiences a ferromagnetic-like collapse.
In Table 1 we summarize the results of varying the shape and magnitude of $\Delta\omega$ (see Appendix A & B). This yields the stability boundaries of the topological gas dynamics and the thermodynamic Onsager attractor.
@@ V Conclusion
We have identified a thermodynamic phase transition in chiral active matter, where intermediate disorder activates what resembles a topological heat pump that drives an inverse energy cascade ($\Pi(k)<0$). This cascade inevitably saturates in a negative-temperature Onsager dipole, a macroscopic ordered state protected by topological invariants. By distinguishing between the microscopic active bath and the macroscopic renormalised fluid, we provide a unified framework that explains the coexistence of seemingly dissipative spectra and inertial transport. These findings suggest that the large-scale organization of active matter may be governed by an inertia that emerges from topological constraints, offering a new paradigm for engineering robust transport in synthetic active fluids.
@@ VI End Matter
The End Matter provides the empirical and theoretical foundation for the phase diagram in Table 1, first through simulations that systematically vary the disorder distribution $\Delta\omega$, then through a rigorous theoretical derivation of the model’s hydrodynamic limit. First, however, we must outline a conjunction of two seemingly disparate turbulence theories.
VI.1 Appendix A: Three thermodynamic phases
@@
Validating Ref. [18]’s concept of negative viscosity, we observe that our simulated active bath pumps energy from the injection scale to the system size, stabilizing the macroscopic state via multiplicative noise. This energy accumulation is remarkably analogous to the self-similar condensate characterized by Ref. [33], where the coherent dipole effectively slaves the background fluctuations to its global topology.
@@
Figure 4: The three distributions in $\Delta\omega$ corresponding to the three thermodynamic phases in Table 1, showing only the positive component ($\omega_{i}$ is symmetric around 0). The lower limit on the $x$-axis is given by $v_{0}/2L$, the hypothetical “box-size fluctuation.”.
Figure 5: Summary of a Phase I simulation (top row) and a Phase II simulation emphasising the emergence of hard-core defects (bottom row). Topological charge is shown with a colorscale; timestamps are indicated.
Figure 6: Ensemble averages of 16 simulations, showing the behaviour of the three phases in Table 1. The leftmost column (panels a, e, and i) show the defect count $N_{d}(t)$ in a solid blue line, with a $t^{-0.75}$ powerlaw in dotted green line and a fitted (through non-linear minimization of root-mean square error) powerlaw in dashed red line. The next column (panels b, f, and j) show median spectral flux for the raw (red) and RFE (blue) fields. The third column (panels c, g, and k) show the energy spectrum for the raw (red) and RFE (blue) fields, with $k^{-3}$ and $k^{-5/3}$ powerlaw fits, as well as the cut-off scale $k_{c}$. The last column (panels d, h, and l) show the thermodynamic trajectory. All quantities are shown as medians of 16 randomized (in initial locations $x_{i},y_{i}$ and white-noise $\eta_{i}$) runs.
@@
To conclude the theoretical conjunction, we show, in Figure 4, three distributions in $\omega_{i}$, corresponding to $\Delta\omega$ in Table 1. A matter of some importance, the intrinsic frustration $\omega_{i}$ near the box-size, $v_{0}/2L$, acts simultaneously as $Z$ in Ref. [18] and $\Omega$ in Ref. [33]. Thus, if the frustration dispersion $\Delta\omega$ is too low and too narrow, the system generates active stress but lacks the transport coherence to move it, yielding an arrested glass vortex (Figure 3, Phase II in Table 1). Reproducing a ferromagnetic-like collapse; a cold, isotropic, uniform “frustration” ($\Delta\omega\rightarrow 0$) yields a collection of static clumps (Phase I in Table 1). Conversely, if the dispersion in $\omega_{i}$ is both sufficiently high and sufficiently broad (near the box-size fluctuations, $v_{0}/2L$), the system effectively facilitates the emergence of a space-filling Onsager dipole (Phase III in Table 1).
In Figure 6 we rigorously demonstrate the foregoing by extracting the median results from 16 ensemble simulations of the three phases. Crucially, we observe that the spectral flux in Phase I (Figure 5) is positive for both the raw and RFE fields, while for Phase II (Figure 6f), the flux is negative for the raw field (crossing zero at intermediate wavenumbers) and positive for the RFE field, and for Phase III the flux is consistently negative, facilitating the inverse cascade that would otherwise be arrested in the case of Phase II.
Thus, by revealing that the hydrodynamic condensate characterized by Ref. [33] is the physical realization of the noise-induced phase transition predicted by Ref. [18], we have provided an independent proof of our central thesis, namely a unified theory of active chiral matter where microscopic phase synchronisation and macroscopic inverse cascades are understood as dual manifestations of the same thermodynamic attractor.
VI.2 Appendix B: Derivation of advective transport

### v4

@@ III Results
The temporal evolution of the system, shown in Figure 1, reveals a spontaneous transition from microscopic chaos ($t=0$) to macroscopic order ($t=11)$, as small-scale phase defects nucleate and annihilate rapidly. Surviving defects begin to cluster and merge, forming a sloshing, large-scale dipole ($t=22$ and beyond).
@@
To discern the mechanism driving this dynamic condensation, we analyzed the kinetic pathways of the topological defects (Figure 2b), the decay of the defect number density $N_{d}(t)$. Adherence to the theoretical $t^{-0.75}$ scaling [38, 40], distinct from the diffusive $t^{-1}$ rate, confirms that vortex merger is governed by dispersive shock dynamics [41, 42, 40] and acoustic horizons [43], rather than passive strain fields or random-walk annihilation [44].
@@
The directionality of the energy transfer is established by the spectral energy flux $\Pi(k)$ (Figure 2c), where the distinct negative plateau in the spectral flux ($\Pi(k)<0$) rigorously proves the inverse directionality of the energy cascade; energy injected at the microscale by $\omega_{i}$ is transported to the box size, creating the Onsager dipole.
@@
A central finding of this work is the resolution of the spectral dichotomy observed in active matter, namely the coexistence of steep, dissipative spectra with large-scale coherent structures [13, 45, 46]. Figure 2d) shows that raw particle velocity field exhibits a steep spectral slope consistent with dissipative enstrophy [33]. This reflects the singular geometry of the topological defects: the raw agents are kinematically pinned to the sharp cusps of the phase field ($\nabla\phi\sim 1/r$), and thus their spectrum measures the packing of vorticity filaments [47].However, when we apply the RFE operator to the ensemble, filtering out the microscopic singularities by averaging over the interaction scale $\tau$, a “hidden” Kolmogorov scaling is revealed: The RFE spectrum $E_{RFE}(k)$ (the blue spectrum in Figure 2d) tracks the classical $k^{-5/3}$ slope characteristic of 2D inviscid fluid turbulence.
@@
Finally, we demonstrate that the formation of the dipole is a thermodynamic inevitability. Figure 2d) maps the system’s trajectory in the energy-order phase plane ($E$ vs. $P$). The system climbs away from the disordered origin (random gas), increasing both its macroscopic kinetic energy and its dipole order parameter $P$, while the flucutations in Figure 2d) reflect the dynamic shock merger mechanism. In the context of Eq. (9), we observe an ascent up the entropy curve into the regime of negative absolute temperature [22]. In the active bath, intrinsic disorder ($\omega_{i}$) continuously injects enstrophy to maximize the entropy of the renormalised fluid, forcing the long-term spectral condensation of vorticity as the only accessible maximum-entropy state within the bounded domain. The negative spectral flux in Figure 2b) confirms our hypothesis that the modeled particles act as a dissipative kinetic layer that drives a superfluid governed by the conservative Kirchhoff-Onsager Hamiltonian, unifying the phenomenology of chiral active matter with the rigorous statistical mechanics of classical inviscid fluids.
@@
Figure 3: Active vortex glass made of particles whose raw distribution features an inverse cascade ($\Pi(k)<0$), but whose RFE field features a forward-cascade ($\Pi(k)>0$). The defect count falls off drastically and flattens; defect merger is arrested and the ensemble settles in a frozen defect lattice. Top row shows the RFE particle point-cloud while bottom row shows topological charge, while the columns correspond to seven temporal snapshots.
@@ IV Discussion
Our results resolve a spectral paradox [48, 49, 4, 50] by demonstrating that chiral active matter may exhibit a fundamental scale-dependent duality. At microscopic scales ($k>k_{c}$), the modeled system is dominated by phase singularities, with a steep spectral slope [33, 34, 35, 36]; in our model, this is not a signature of dissipation, but of enstrophy injection; that is, the conversion of intrinsic chemical energy ($\omega_{i}$) into topological charge.
@@
I. Global synchronisation $\Delta\omega\to 0$ $\nabla\phi\to 0$. Static clumps (Fig. 5). $\Pi_{Raw}>0,\;\Pi_{RFE}>0$
II. Active Vortex Glass Low / narrow Frustration; no merger. Defect solid (Fig. 3). $\Pi_{Raw}<0,\;\Pi_{RFE}>0$
III. Onsager Condensate High / broad Shock merger. Onsager dipole (Fig. 1). $\Pi_{Raw}<0,\;\Pi_{RFE}<0$
At macroscopic scales ($k<k_{c}$), the system reveals a renormalised fluid. By filtering out the singular defects via the RFE operator (Eqs. 4, 5 in empirical terms, Eq. 12 in theoretical terms), we recover an inverse cascade (Figure 2b). The collective motion of the defects generates a “phantom inertia,” an effective advective term, $R(\mathbf{v}\cdot\nabla)\mathbf{v}$ producing dynamics that are isomorphic with shallow water theory (see the Supplementary Materials), despite the overdamped nature of the individual agents.
@@
Crucially, the RFE framework suggests a general resolution for the non-universal spectra frequently reported in biological and active matter experiments [10]. We propose that the sharp deviations from Kolmogorov scaling observed in bacterial swarms, active nematics, and cytoskeletal gels [1, 10, 11, 7, 3, 51] may be artifacts of analyzing the raw particle fields rather than the renormalised hydrodynamic modes. By applying the RFE operator (Eqs. 4, 5), essentially a low-pass filter, it may be possible to recover inertial scaling laws in a broad class of overdamped systems.
Supersonic topological gas dynamics
The realization of the Onsager dipole depends strictly on the system’s ability to overcome topological energy barriers [25]. As alluded to by Ref. [21], we find that the intrinsic frequency dispersion $\Delta\omega$ determines whether the topological gas freezes, jams, or flows.
Figure 3 illustrates the foregoing by presenting the evolution of a simulation that used a narrower, lower-frequency dispersion $\Delta\omega$ (see Figure 4 in the End Matter); the raw particle distribution subsequently drove an inverse cascade, but the renormalised fluid did not, and the active bath was unable to transfer energy to wavenumbers $k<k_{c}$, resulting in an arrested toplogical defect glass state. While the Onsager fluid exhibits continuous, power-law coarsening, the vortex glass state exhibits a rapid localized quench ($N\propto t^{-2}$) followed by a kinetic arrest ($N\rightarrow\text{const.}$), and static, quantized loop currents of phase information [21, 29]. This leads us to propose that the experimental ’active glass’ described in Ref. [52] may be an intrinsic phase of topological gases, for which we provide an explicit theoretical foundation.
In Table 1 we summarize the results of varying the shape of $\Delta\omega$ (see the End Matter). This yields the stability boundaries of the topological gas dynamics and the thermodynamic Onsager attractor, where we identify the limit $R\rightarrow 1$, leading to a ferromagnetic-like collapse.
The Onsager dipole state, which represents the hydrodynamic limit, is thus in a marginal synchronisation regime, meaning that the synchronising force is balanced by continous, shock-driven thermalization of inertial energy. This is empirically confirmed by observing that the dynamics in Figure 1 effectively never stop. The observed density decay $N(t)\sim t^{-0.75}$ during this phase quantitatively distinguishes the regime from compressible gas dynamics ($t^{-2/3}$ or $t^{-1}$, [53]), favoring coherent structure merger in shallow waters [40]. The dominance of shock-mergers over smooth thinning confirms the system operates in the supersonic regime ($M>1$), strictly isomorphic to supercritical shallow water flow, where density fluctuations map directly to hydraulic jumps [43, 41].
@@ V Conclusion
We have identified a disorder-induced transition in chiral active matter where a topological heat pump drives an inverse cascade, saturating in a negative-temperature, dynamic Onsager dipole, representing a hydrodynamic limit that is isomorphic with shallow water theory. This mapping reconciles dissipative microphysics with macroscopic inertial transport. Our findings suggest active matter is governed by an emergent topological inertia, offering a new paradigm for robust transport.
@@ VI End Matter
The End Matter provides the empirical and theoretical foundation for the phase diagram in Table 1, through simulations that systematically vary the disorder distribution $\Delta\omega$, and concludes with some perspectives on the conceptual implications of our results.
The three thermodynamic phases of topological gas dynamics
@@
Figure 4: The three distributions in $\Delta\omega$ corresponding to the three thermodynamic phases in Table 1, showing only the positive component ($\omega_{i}$ is symmetric around 0). The lower limit on the $x$-axis is given by $(2\Lambda)^{-1}$, the hypothetical “box-size fluctuation.”.
@@
Before we pinpoint the conjunction between the latter two theoretical forays, we show, in Figure 4, three distributions in $\omega_{i}$, corresponding to $\Delta\omega$ in Table 1. Intrinsic frustration $\omega_{i}$ near the box-size, $(2\Lambda)^{-1}$, acts simultaneously as $Z$ in Ref. [25] and $\Omega$ in Ref. [36]. Thus, if frustration $\Delta\omega$ is too low-valued and too narrow, the system generates active stress but lacks the transport coherence to move it, yielding an arrested glass vortex. Conversely, if the dispersion in $\omega_{i}$ is both sufficiently high and sufficiently broad (reaching near-box-size fluctuations), the system effectively facilitates the emergence of a space-filling Onsager dipole. In Figure 6 we rigorously demonstrate the foregoing descriptions by extracting the median results from 16 ensemble simulations of the three phases. Crucially, we observe that the spectral flux in Phase I (see Figure 5) is positive for both the raw and RFE fields, while for Phase II (Figure 6f), the flux is negative for the raw field (crossing zero at intermediate wavenumbers) and positive for the RFE field, and for Phase III the flux is consistently negative, facilitating the inverse cascade that would otherwise be arrested in the case of Phase II.
Figure 5: Summary of a Phase I simulation.. RFE particles as point-clouds (top row) and topological charge (bottom row) mapped with a colorscale; timestamps are indicated.
Figure 6: Ensemble averages of 16 simulations, showing the behaviour of the three phases in Table 1. The leftmost column (panels a, e, and i) show the defect count $N_{d}(t)$ in a solid blue line, with a $t^{-0.75}$ powerlaw in dotted green line and a fit for $t>3.4$ (through non-linear minimization of root-mean square error) powerlaw in dashed red line. The next column (panels b, f, and j) show median spectral flux for the raw (red) and RFE (blue) fields. The third column (panels c, g, and k) show the energy spectrum for the raw (red) and RFE (blue) fields, with $k^{-3}$ and $k^{-5/3}$ powerlaw fits, as well as the cut-off scale $k_{c}$. The last column (panels d, h, and l) show the thermodynamic trajectory. All quantities are shown as medians of 16 randomized (in initial locations $x_{i},y_{i}$ and white-noise $\eta_{i}$) runs.
Thus, the effective Euler dynamics we observe allow the system to access the condensate state. This suggests that Ref. [25]’s thermodynamic driver is strong enough in this regime to overcome the chaotic transients [16] and access Ref. [36]’s equilibrium basin. This is in contrast with the active nematic systems that often saturate in a state of defect chaos [16]. Thus, the theoretical conjunction between Ref. [25]’s extremum principles with Ref. [36]’s statistical mechanics bolsters the notion that the Onsager dipole in our simulation is a thermodynamically preferred attractor.
Future studies should map out the biological, chemical, and environmental circumstances that facilitate the emergence of an inertial flow in soft, active matter turbulence, and when such systems are dominated by low-Reynolds number friction and rapid energy decay [14, 15, 16], as well as specific instances where the theoretical foundations for our hydrodynamic limit breaks down.
The potential reach of supersonic topological gas dynamics
In the biological microscopic realm, large-scale mixing is vital for nutrient transport [54]. Conceptually, our model suggests that a biological system benefits from a specific variance in intrinsic traits (modeled as the dispersion $\Delta\omega$) to activate the topological heat pump, or revert to a defect lattice. This provides a mechanistic explanation for why phenotypic heterogeneity is prevalent [55] and suggests that such noise expressions are in fact beneficial for cohesion. Biological systems may thus inadvertently exploit noise in gene expression [56], effectively tapping into disorder as a means for macroscopic transport.
In our model, when the system is disordered ($R\to 0$), it is effectively massless and governed by overdamped diffusion. As it synchronizes, it acquires mass ($\lambda=R^{2}$), supporting momentum transport and waves akin to correlation propagation in flocks and swarms [14, 57], with the caveat that our system operates in two dimensions. The observed $R$-dependency implies that defects behave as massless holes (see Figure 7 in the Supplementary Materials) in an otherwise massive fluid. This mass contrast enables a supersonic topological gas dynamic: defects accelerate across phase gradients faster than the bulk sound speed [58], maintaining scale-free coherence [59] without central processing.
In the field of robotics, heterogeneity is usually considered a manufacturing error. For example, the Kuramoto model describes the superconductor-to-insulator transition (or Bose Metal state) in dirty Josephson arrays, quantum oscillator systems that feature synchronisation due to fabrication imperfections [60, 24], to which the transition between Phases II and III in Table 1 is (classically and locally) analogous. This heterogeneity, which typically cannot be avoided, can potentially be exploited as a control input [61, 44]. For example, in smart materials, the Phase II-III transition can be achieved by electromagnetic field-control of rotors, enabling shock-absorption (converting kinetic energy temporarily into topological sorting; see also Refs. [62, 63]). With attention to edge-effects, microfluidic applications could involve an “Onsager valve” with zero moving parts and negative effective viscosity, acting both as valve and pump, depending on control input (see Ref. [64]).
VII Supplementary Materials: Derivation of the hydrodynamic limit